# Unveiling the Blind Spots: A Critical Examination of Fairness in Autonomous Driving Systems

### v2

@@ Abstract.
Autonomous driving systems have extended the spectrum of Web of Things for intelligent vehicles and have become an important component of the Web ecosystem. Similar to traditional Web-based applications, fairness is an essential aspect for ensuring the high quality of autonomous driving systems, particularly in the context of pedestrian detectors within them. However, there is an absence in the literature of a comprehensive assessment of the fairness of current Deep Learning (DL)-based pedestrian detectors. To fill the gap, we evaluate eight widely-explored DL-based pedestrian detectors across demographic groups on large-scale real-world datasets. To enable a thorough fairness evaluation, we provide extensive annotations for the datasets, resulting in 8,311 images with 16,070 gender labels, 20,115 age labels, and 3,513 skin tone labels. Our findings reveal significant fairness issues related to age. The undetected proportions for adults are 20.14% lower compared to children. Furthermore, we explore how various driving scenarios affect the fairness of pedestrian detectors. We find that the bias may exacerbate for children and females towards low brightness and low contrast.
copyright: nonecopyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 1. Introduction
Although significant testing efforts have been made, to the best of our knowledge, the study of fairness testing for autonomous driving systems remains under-investigated in the literature. From a Software Engineering (SE) perspective, fairness is considered a non-functional software property, making it an important subject for testing (Zhang et al., 2022). Fairness testing, as an emerging domain within software testing, seeks to uncover fairness issues in software systems (Galhotra et al., 2017).
Fairness issues in autonomous driving systems, such as a higher accuracy in detecting pedestrians of white ethnicity compared to black ethnicity, can perpetuate discriminatory outcomes and unequal treatment based on race. This can result in harm to individuals belonging to marginalized groups, further exacerbating existing social inequalities. Therefore, it is crucial to prioritize fairness testing in autonomous driving systems.
To fill the gap, we investigate fairness of eight state-of-the-art DL-based pedestrian detectors that have been extensively studied in the research community. Our main focus is to quantitatively assess performance disparities in these detectors across diverse demographic groups, which are widely recognized as group fairness issues (Chen et al., 2022b). To enable this investigation, we manually enrich four widely-adopted real-world datasets with gender, age, and skin tone labels, resulting in a collection of 8,311 real-world images annotated with 16,070 gender labels, 20,115 age labels, and 3,513 skin tone labels. Using these labeled datasets, we assess the group fairness of existing pedestrian detectors and also explore how commonly-studied driving scenarios (including various brightness, contrast, and weather conditions) impact the fairness of these detectors.
Our study reveals the following findings: (1) Overall, state-of-the-art pedestrian detectors exhibit significant bias across different age and skin tone groups. On the four datasets examined, the undetected proportions for children surpass those for adults by an average of 20.14%. (2) However, the overall performance of these pedestrian detectors in detecting males and females and dark-skin and light-skin groups does not exhibit a large difference, with only a 1.19% and 0.44% gap in undetected proportions. (3) Regarding the exploration of common scenarios, the detection bias towards the children and females increases under low-brightness and low-contrast conditions compared to adults and males. For instance, the difference in undetected proportions for adults and children increases from 22.45% to 25.48% between day time and night time scenarios.
@@ 2.1. Software Fairness
Fairness has gained significant attention in the SE community since its initial exploration by SE researchers in 2008 (Finkelstein et al., 2008). There have been various definitions of fairness in the literature. In this paper, we focus on group fairness, a concept extensively studied in software fairness research (Biswas and Rajan, 2021; Zhang and Harman, 2021; Hort et al., 2021; Chakraborty et al., 2021, 2020; Gohar et al., 2023; Biswas and Rajan, 2020; Zhang et al., 2023; Zhang and Sun, 2022). Notably, group fairness closely aligns with legal regulations on fairness (Barocas and Selbst, 2016), such as the adherence to the four-fifths rule, a cornerstone of US law (Wick et al., 2019). Consequently, prioritizing group fairness when building software has emerged as an essential ethical duty and requirement for software engineers (Chakraborty et al., 2021).
In the context of group fairness, certain personal characteristics that require protection against unfairness during decision-making are called sensitive attributes, also known as protected attributes (Zhang et al., 2022; Corbett-Davies and Goel, 2018). Well-recognized sensitive attributes include race, sex, age, pregnancy, familial status, disability status, and more (Chen et al., 2022a). These sensitive attributes typically partition a population into distinct groups: a privileged group and an unprivileged group (Chen et al., 2022a). Group fairness entails the equal treatment of these groups by the same machine learning model. However, in practice, members of unprivileged groups often experience systematic disadvantages, resulting from unfair machine learning models. For instance, in the context of a pedestrian detection task, if age is deemed a sensitive attribute, the predictive model may exhibit a bias favoring adult groups over child groups. In this scenario, the adult group is considered the privileged group, while the child group becomes the unprivileged one.
Presently, the majority of software fairness testing research primarily revolves around tabular data, as indicated by a recent survey (Chen et al., 2022a). This paper centers on fairness testing for pedestrian detection in autonomous driving systems. We specifically examine three sensitive attributes (i.e., gender, age, and skin tone) that are more recognizable in autonomous driving datasets and have been widely considered in the fairness literature (Soremekun et al., 2022).
@@ 2.2. Autonomous Driving Testing
Autonomous driving is a hot SE research topic, and researchers have proposed various testing techniques for autonomous driving systems (Pei et al., 2019; Tian et al., 2018; Zhang et al., 2018; Zhou et al., 2020; Guo et al., 2022). These techniques mainly involve the automatic generation of artificial driving scenarios to simulate real-world conditions, enabling the evaluation of system behavior and the detection of blind spots or corner bugs across various autonomous driving scenarios.
While a substantial body of knowledge focuses on assessing the robustness and correctness properties of autonomous driving systems (Zhang et al., 2022), to the best of our knowledge, only few studies have explored the fairness properties, particularly in the pedestrian detection domain within autonomous driving. Pedestrian detection is a crucial process identifies pedestrians within street-level images by providing their predicted locations along with corresponding bounding boxes and confidence scores (Dollár et al., 2012; Zhang et al., 2016). Brandao (Brandao, 2019) discussed fairness in pedestrian detection but focused solely on traditional detection techniques, which are no longer dominant in contemporary practice. Wilson et al. (Wilson et al., 2019) examined skin tone bias but with limitations to one dataset. Similarly, Kogure et al. (Kogure et al., 2022) only explored age bias with one detection method and a small-scale dataset, lacking a comprehensive analysis. To summarise, current fairness research in pedestrian detection of autonomous driving is still in its early stages and faces limitations in terms of diversity in pedestrian detectors, testing datasets, and sensitive attributes considered.
To address this knowledge gap, our paper presents the a systematic empirical study on group fairness testing in pedestrian detection. We conduct experiments using eight popular detection methods and four diverse testing datasets, encompassing different scenarios determined by a variety of factors such as brightness, contrast, and weather conditions. Through this comprehensive study, we gain insights into the potential bias and risks associated with using existing pedestrian detectors.
@@ 3. Experimental Design
This section introduces our research questions and experimental settings.
@@ 3.1. Research Questions
RQ2 (Fairness in different scenarios): What fairness do pedestrian detectors achieve in different brightness, contrast, and weather conditions? We further investigate the fairness of commonly-studied pedestrian detectors in different autonomous driving scenarios by considering a variety of different brightness (RQ2.1), contrast (RQ2.2), and weather conditions (RQ2.3) (Tian et al., 2018; Zhang et al., 2018; Zhou et al., 2020).
In the following, we introduce the pedestrian detectors (Section 3.2), benchmark datasets (Section 3.3), evaluation metrics (Section 3.4), and statistical analysis (Section 3.5) that we use to address the above research questions.
3.2. Pedestrian Detection Methods
In recent years, DL has revolutionized pedestrian detection approaches. We focus our analysis on eight DL-based pedestrian detectors that are widely studied in the autonomous driving community (Zou et al., 2023; Cao et al., 2020). These detectors are pre-trained DL models that researchers and practitioners can directly use for pedestrian detection tasks. They can be classified into two categories (Hasan et al., 2020): general object detectors and pedestrian-specific detectors. Next, we briefly introduce each category and its pedestrian detectors we study. Table 1 provides an overview of these detectors (“Detector” shows the name of a pedestrian detector; “Backbone/Model” represents the architecture of the detector, which could be either the specific variant of the model employed, such as ‘YOLOX-Tiny’, ‘YOLOX-S’, ‘YOLOX-L’, or ‘YOLOX-X’ in the case of the YOLOX series, or the pre-trained deep neural network used for extracting features from input images or the model size we choose; “Source” indicates the framework/toolkit name and its source for a given pedestrian detector).
@@ 3.2. Pedestrian Detection Methods
Type Detector Backbone/Model Source
General object detectors YOLOX YOLOX-X MMDetection (MMDetection Developers, nd)
@@
General object detectors: General detectors can detect various objects such as cars, traffic lights, and pedestrians. They have great generalization ability but lack pedestrian-specific adaptation (Hasan et al., 2020). They can be categorized into two genres (Zou et al., 2023): two-stage and one-stage detectors. Two-stage detectors propose regions before feature extraction and classification, achieving high accuracy but slower speed; one-stage detectors complete all operations in one step, providing faster speed but lower accuracy. Both types involve trade-offs and are widely used for pedestrian detection. Hence, this paper selects detectors from both categories. For one-stage detectors, we adopt the widely-studied YOLOX (Ge et al., 2021) (a faster extension of the YOLO series (Redmon et al., 2016)) and RetinaNet (Lin et al., 2017) (which addresses the class imbalance problem). For two-stage detectors, we employ the Faster R-CNN (Ren et al., 2015) (one of the pioneering detectors in the R-CNN family) and Cascade R-CNN (Cai and Vasconcelos, 2018) (which achieves higher accuracy through a cascade of multiple CNNs to refine region proposals). These detectors have been extensively used in the autonomous driving literature (Hasan et al., 2020; Cao et al., 2020; Yu et al., 2020; Braun et al., 2019).
@@ 3.3.2. Sensitive Attribute Labeling
As introduced in Section 2.1, we focus on three sensitive attributes (i.e., gender, age, and skin tone) of pedestrians that are recognizable in autonomous driving images and widely studied in the fairness literature (Soremekun et al., 2022). To enable fairness testing, we need datasets with labels that identify these sensitive attributes of the humans depicted in the images. Among the datasets investigated herein, the only sensitive attribute already labeled is the skin tone (i.e., light-skin tone and dark-skin tone) for the BDD100k dataset. Therefore, we manually label the other sensitive attributes for each of the datasets considered in this study.
@@
Gender Age Skin
Dataset Male Female Adult Child LS DS
@@ 3.3.3. Scenario Labeling
Labeling weather conditions. Common weather conditions studied in the autonomous driving literature include rain, fog, and snow (Tian et al., 2018; Zhou et al., 2020; Zhang et al., 2018; Pei et al., 2019). However, our datasets rarely contain images depicting fog and snow. This is due to the fact that our datasets are collected from real-world scenarios where fog and snow are infrequently encountered. The limited samples of snowy or foggy weather pose challenges for statistical analysis. As a result, we focus on rain as the weather condition of interest. Two annotators independently classify images containing labeled pedestrians into two categories: rainy and non-rainy. During this process, 1,856 images are not annotated because the two annotators and the arbitrator all cannot accurately distinguish the weather conditions.

### v3

# Bias Behind the Wheel: Fairness Analysis of Autonomous Driving Systems

@@ Abstract.
This paper analyzes fairness in automated pedestrian detection, a crucial but under-explored issue in autonomous driving systems. We evaluate eight state-of-the-art deep learning-based pedestrian detectors across demographic groups on large-scale real-world datasets. To enable thorough fairness testing, we provide extensive annotations for the datasets, resulting in 8,311 images with 16,070 gender labels, 20,115 age labels, and 3,513 skin tone labels. Our findings reveal significant fairness issues, particularly related to age. The undetected proportions for children are 20.14% higher compared to adults. Furthermore, we explore how various driving scenarios affect the fairness of pedestrian detectors. We find that pedestrian detectors demonstrate significant gender biases during night time, potentially exacerbating the prevalent societal issue of female safety concerns during nighttime out. Moreover, we observe that pedestrian detectors can demonstrate both enhanced fairness and superior performance under specific driving conditions, which challenges the fairness-performance trade-off theory widely acknowledged in the fairness literature. We publicly release the code, data, and results to support future research on fairness in autonomous driving.
Fairness, pedestrian detection, autonomous driving
copyright: acmcopyrightjournal: TOSEMjournalyear: 2024journalvolume: 1journalnumber: 1article: 1publicationmonth: 7price: 15.00price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software creation and managementccs: Computing methodologies Machine learning
@@ 1. Introduction
Although significant testing efforts have been made, to the best of our knowledge, the study of fairness testing for autonomous driving systems remains under investigation in the literature. From a Software Engineering (SE) perspective, fairness is considered a non-functional software property, making it an important subject for testing (Zhang et al., 2022). Fairness testing, as an emerging domain within software testing, seeks to uncover fairness issues in software systems (Chen et al., 2024).
Fairness issues in autonomous driving systems, such as a higher accuracy in detecting male pedestrians compared to females, can perpetuate discriminatory outcomes and unequal treatment based on gender. This can result in harm to individuals belonging to marginalized groups, further exacerbating existing social inequalities. Therefore, it is crucial to prioritize fairness testing in autonomous driving systems.
To fill the gap, we investigate fairness of eight state-of-the-art Deep Learning (DL)-based pedestrian detectors that have been extensively studied in the research community. Our main focus is to quantitatively assess performance disparities in these detectors across diverse demographic groups, which are widely recognized as group fairness issues (Chen et al., 2022). To enable this investigation, we manually enrich four widely-adopted real-world datasets with gender, age, and skin tone labels, resulting in a collection of 8,311 real-world images annotated with 16,070 gender labels, 20,115 age labels, and 3,513 skin tone labels. Using these labeled datasets, we assess the group fairness of existing pedestrian detectors and also explore how commonly-studied driving scenarios (including various brightness, contrast, and weather conditions) impact the fairness of these detectors.
Our study reveals the following findings: (1) Overall, state-of-the-art pedestrian detectors exhibit significant bias regarding age. On the four datasets examined, the undetected proportions for children surpass those for adults by an average of 20.14%. However, the overall performance of these pedestrian detectors in detecting males and females and dark-skin and light-skin groups does not exhibit a large difference, with only a 1.19% and 0.44% gap in undetected proportions. (2) The studied pedestrian detectors reveal significant gender biases during night time, with a higher proportion of females going undetected compared to males. This situation may aggravate existing societal concerns about female safety during nighttime outings. (3) In contrast to the commonly accepted fairness-performance trade-off, our findings suggest that pedestrian detectors can achieve enhanced fairness and detection performance under specific driving scenarios, such as those with higher brightness levels.
@@ 2.1. Software Fairness
Fairness has gained significant attention in the SE community since its initial exploration by SE researchers in 2008 (Finkelstein et al., 2008). There have been various definitions of fairness in the literature. In this paper, we focus on group fairness, a concept extensively studied in software fairness research (Biswas and Rajan, 2021; Zhang and Harman, 2021; Hort et al., 2021; Chakraborty et al., 2021, 2020; Gohar et al., 2023; Biswas and Rajan, 2020a; Zhang et al., 2023; Zhang and Sun, 2022; Nguyen et al., 2023). Notably, group fairness closely aligns with legal regulations on fairness (Barocas and Selbst, 2016), such as the adherence to the four-fifths rule, a cornerstone of US anti-discrimination law (Biddle, 2005; Feldman et al., 2015; Dwork et al., 2012; Zafar et al., 2017; Wick et al., 2019). Consequently, testing and prioritizing group fairness when building software has emerged as an essential ethical duty and requirement for software engineers (Chakraborty et al., 2021).
In the context of group fairness, certain personal characteristics that require protection against unfairness during decision-making are called sensitive attributes, also known as protected attributes (Zhang et al., 2022; Corbett-Davies and Goel, 2018). Well-recognized sensitive attributes include race, sex, age, pregnancy, familial status, disability status, and more (Chen et al., 2024; Wan et al., 2023). These sensitive attributes typically partition a population into distinct groups: a privileged group and an unprivileged group (Chen et al., 2024). Group fairness entails the equal treatment of these groups by the same machine learning model. However, in practice, members of unprivileged groups often experience systematic disadvantages, resulting from unfair machine learning models. For instance, in the context of a pedestrian detection task, if age is deemed a sensitive attribute, the predictive model may exhibit a bias favoring adult groups over child groups. In this scenario, the adult group is considered the privileged group, while the child group becomes the unprivileged one.
Recently, Chen et al. (Chen et al., 2024) has presented a comprehensive survey of fairness testing research and analyzed its trend. This survey points out that the majority of existing work revolves around tabular data (Biswas and Rajan, 2020b, 2021; Gohar et al., 2023; Zhang and Harman, 2021; Chen et al., 2023). For example, Biswas and Rajan (Biswas and Rajan, 2020b) evaluated fairness of machine learning models on a crowd sourced platform using tabular datasets covering tasks such as credit risk prediction, income prediction, marketing, and loan application. Similarly, Chen et al. (Chen et al., 2023) conducted an empirical study on the group fairness achieved by state-of-the-art bias mitigation methods across eight commonly used tabular data-driven decision tasks. In the contrast, our paper centers on fairness testing for pedestrian detection in autonomous driving systems. We specifically examine three sensitive attributes (i.e., gender, age, and skin tone) that are recognizable in autonomous driving datasets. These sensitive attributes have been demonstrated to be the most widely considered ones in the fairness testing literature (Chen et al., 2024).
@@ 2.2. Autonomous Driving Testing
Autonomous driving is a hot SE research topic, and researchers have proposed various testing techniques for autonomous driving systems (Pei et al., 2017; Tian et al., 2018; Zhang et al., 2018; Zhou et al., 2020; Guo et al., 2022; Zuo et al., 2021). For instance, Tian et al. (Tian et al., 2018) proposed DeepTest, a novel technique using image transformations to emulate potential camera disturbances encountered in driving environments. Zhang et al.  (Zhang et al., 2018) presented DeepRoad, employing Generative Adversarial Networks to craft test images derived from actual driving scenarios. Zhou et al.  (Zhou et al., 2020) introduced DeepBillboard, aiming to generate adversarial billboards that could lead to steering mistakes in autonomous vehicles. Guo et al. (Guo et al., 2022) developed LiRTest, marking the inaugural automated testing technique for LiDAR-equipped autonomous driving software.
While a substantial body of knowledge focuses on assessing the robustness and correctness properties of autonomous driving systems (Zhang et al., 2022), to the best of our knowledge, only a few studies have explored the fairness properties, particularly in the pedestrian detection domain within autonomous driving, indicating that this area remains underdeveloped.
Pedestrian detection is a crucial process that identifies pedestrians within street-level images by providing their predicted locations along with corresponding bounding boxes and confidence scores  (Dollár et al., 2012; Zhang et al., 2016). Despite its significance, research on the fairness of pedestrian detection within autonomous driving is limited. Brandao (Brandao, 2019) explored fairness in pedestrian detection, concentrating only on classic machine learning-based methods. These classic techniques, reliant on manually-defined features, have been eclipsed by deep learning-based pedestrian detection approaches, now prevalent in the autonomous driving industry. Wilson et al.  (Wilson et al., 2019) focused on skin tone bias, confined their analysis to a single dataset and two general object detectors. Similarly, Kogure et al.  (Kogure et al., 2022) explored only age bias using a small-scale dataset and a detection method that is no longer state-of-the-art.
In summary, current fairness studies in pedestrian detection suffer from a lack of variety in the pedestrian detectors evaluated, the datasets used, and the range of sensitive attributes explored. Furthermore, no previous work has explored how different environmental characteristics (e.g., brightness, contrast, and weather conditions) affect fairness.
To address this knowledge gap, our paper presents a comprehensive empirical study on revealing fairness issues in pedestrian detection. We conduct experiments using eight popular detection methods and four diverse testing datasets, encompassing different scenarios determined by a variety of factors such as brightness, contrast, and weather conditions. The scale and diversity of our experiments enable us to provide comprehensive insights into the fairness circumstances of the existing pedestrian detectors.
@@ 3. Experimental Design
This section introduces our research questions and experimental design.
Figure 1. Overview of our experimental settings.
@@ 3.1. Research Questions
RQ2 (Fairness in different scenarios): What fairness do pedestrian detectors achieve in different brightness, contrast, and weather conditions? We further investigate the fairness of commonly-studied pedestrian detectors in different autonomous driving scenarios by considering a variety of brightness (RQ2.1), contrast (RQ2.2), and weather conditions (RQ2.3) (Tian et al., 2018; Zhang et al., 2018; Zhou et al., 2020).
Figure 1 illustrates our experimental settings to answer these RQs. In the following, we introduce the pedestrian detectors, benchmark datasets, evaluation metrics, and experimental details.
3.2. Pedestrian Detectors
In recent years, DL has revolutionized pedestrian detection approaches. We focus our analysis on eight DL-based pedestrian detectors that are widely studied in the autonomous driving community (Zou et al., 2023; Cao et al., 2020). These detectors are pre-trained DL models that researchers and practitioners can directly use for pedestrian detection tasks. They can be classified into two categories (Hasan et al., 2020): general object detectors and pedestrian-specific detectors. Next, we briefly introduce each category and the corresponding pedestrian detectors.
Table 1 provides an overview of these detectors. “Detector” shows the name of a pedestrian detector; “Backbone” represents the pre-trained deep neural network used for extracting features from input images; “Source” indicates the framework/toolkit name and its source for a given pedestrian detector.
@@ 3.2. Pedestrian Detectors
Type Detector Backbone Source
General object detectors YOLOX - MMDetection (MMDetection Developers, nd)
@@
General object detectors: General detectors can detect various objects such as cars, traffic lights, and pedestrians. They have great generalization ability but lack pedestrian-specific adaptation (Hasan et al., 2020). They can be categorized into two genres (Zou et al., 2023): two-stage and one-stage detectors. Two-stage detectors propose regions before feature extraction and classification, achieving high accuracy but slower speed; one-stage detectors complete all operations in one step, providing faster speed but lower accuracy. Both types involve trade-offs and are widely used for pedestrian detection. Hence, this paper selects detectors from both categories. For one-stage detectors, we adopt the widely-studied YOLOX (Ge et al., 2021) (a faster extension of the YOLO series (Redmon et al., 2016) used in Apollo’s autonomous driving systems (Apo, 2023)) and RetinaNet (Lin et al., 2017) (which addresses the class imbalance problem). For two-stage detectors, we employ the Faster R-CNN (Ren et al., 2017) (one of the pioneering detectors in the R-CNN family) and Cascade R-CNN (Cai and Vasconcelos, 2018) (which achieves higher accuracy through a cascade of multiple CNNs to refine region proposals). These detectors have been extensively used in the autonomous driving literature (Hasan et al., 2020; Cao et al., 2020; Yu et al., 2020; Braun et al., 2019).
@@ 3.3.2. Sensitive Attribute Labeling
We focus on three sensitive attribute: gender, age, and skin tone. These attributes are identifiable in autonomous driving images and are recognized as the three most extensively studied sensitive attributes in fairness testing literature (Chen et al., 2024). To enable fairness analysis, we need datasets with labels that indicate these sensitive attributes of the humans depicted in the images. Among the datasets investigated herein, the only sensitive attribute already labeled is the skin tone (i.e., light-skin tone and dark-skin tone) for the BDD100k dataset.
We manually label gender and age for each of the datasets considered in this study. For skin tone, we only use the BDD100k dataset. As described in Section 3.3.1, this dataset shows a greater diversity in skin tones compared to other datasets. This diversity is linked to the geographic locations where the datasets were collected. Specifically, BDD100k, collected in the United States, exhibits more varied skin tones. In contrast, other datasets from European cities, such as CityPersons collected in Germany, have significantly fewer dark-skin individuals. As reported (pew, 2024; wik, 2024a, b), there is only around 1% representation of individuals with dark skin in Germany. During our labeling process, we indeed encountered scarcity of dark-skin individuals in the CityPersons, EuroCity-Day, and EuroCity-Night datasets, leading to our decision not to label these three datasets with skin tone attributes.
@@
Gender Age Skin tone
Dataset Male Female Adult Child Light Dark
@@ 3.3.3. Scenario Labeling
Labeling weather conditions. Common weather conditions studied in the autonomous driving literature include rain, fog, and snow (Tian et al., 2018; Zhou et al., 2020; Zhang et al., 2018; Pei et al., 2017). However, our datasets rarely contain images depicting fog and snow. This is due to the fact that our datasets are collected from real-world scenarios where fog and snow are infrequently encountered. The limited samples of snowy or foggy weather pose challenges for statistical analysis. As a result, we focus on rain as the weather condition of interest. Two annotators independently classify images containing labeled pedestrians into two categories: rainy and non-rainy. During this process, 1,856 images are not annotated because the two annotators and the arbitrator all cannot accurately distinguish the weather conditions.